OpenStudy (anonymous):

How do I write two eaquations in standard form that are equivalent to the given equation: 8x-14y=2

OpenStudy (anonymous):

Multiply 8x-14y=2 by 2 ( 8x-14y=2)*2= 16x-28y=4

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

It's asking for standard form, not y-intercept form. Therefore y=-4/7x+-1/7 would't work.

OpenStudy (anonymous):

Standard form: Ax+By=C y-intercept form: y=mx+b

OpenStudy (anonymous):

thats right sorry can you answer my question plz